### v1

# Recurrent neural network wave functions for Rydberg atom arrays on kagome lattice

Abstract
Rydberg atom array experiments have demonstrated the ability to act as powerful quantum simulators, preparing strongly-correlated phases of matter which are challenging to study for conventional computer simulations. A key direction has been the implementation of interactions on frustrated geometries, in an effort to prepare exotic many-body states such as spin liquids and glasses. In this paper, we apply two-dimensional recurrent neural network (RNN) wave functions to study the ground states of Rydberg atom arrays on the kagome lattice. We implement an annealing scheme to find the RNN variational parameters in regions of the phase diagram where exotic phases may occur, corresponding to rough optimization landscapes. For Rydberg atom array Hamiltonians studied previously on the kagome lattice, our RNN ground states show no evidence of exotic spin liquid or emergent glassy behavior. In the latter case, we argue that the presence of a non-zero Edwards-Anderson order parameter is an artifact of the long autocorrelations times experienced with quantum Monte Carlo simulations. This result emphasizes the utility of autoregressive models, such as RNNs, to explore Rydberg atom array physics on frustrated lattices and beyond.
I Introduction
Rydberg atom arrays have emerged as a rich playground for quantum simulation of many-body problems Browaeys and Lahaye (2020). A key property of these arrays is their high degree of programmability, which enables the realization of multiple Hamiltonians on different lattice geometries and parameter ranges. This programmability facilitates the simulation of a wide array of phases of matter Ebadi et al. (2021); Wurtz et al. (2023) and enables the solution to challenging combinatorial optimization problems Wurtz et al. (2023); Ebadi et al. (2022); Nguyen et al. (2023). Remarkably, the preparation of spin liquid phases—disordered phases of matter characterized by the presence of anyonic excitations, topological invariants, and long-range entanglement—has been demonstrated in programmable Rydberg arrays, potentially serving as building blocks of future generation of fault-tolerant qubits Dennis et al. (2002); Kitaev (2006); Kitaev and Laumann (2009).
Recent numerical studies have investigated the physics of the ground state of Rydberg atom arrays in different lattice geometries, in particular in one Samajdar et al. (2018) and two spatial dimensions in various geometries Samajdar et al. (2020); Kalinowski et al. (2022); Li et al. (2022); Samajdar et al. (2021); Verresen et al. (2021); Yang and Xu (2022); Kornjača et al. (2023). In lattices such as ruby and honeycomb lattices, strong numerical evidence favours the existence of a spin liquid phase in agreement with experiments Verresen et al. (2021); Kornjača et al. (2023). Another recent example is the kagome lattice, where Density Matrix Renormalization Group (DMRG) White (1992); Schollwöck (2011) studies provided evidence that Rydberg atom arrays host a liquid-like regime Samajdar et al. (2021), while Quantum Monte Carlo (QMC) simulations predicted the existence of a spin glass phase Yan et al. (2023). These systems display frustration arising from lattice geometry and Hamiltonian interactions, leading to the existence of a large number of quantum states with nearly degenerate energies but markedly different properties. This makes it computationally difficult to accurately approximate the ground state of these systems.
Here we focus on applying recurrent neural network (RNNs) wave functions  Hibat-Allah et al. (2020); Roth (2020) to a Rydberg array of atoms on the kagome lattice. The effectiveness of RNNs and Transformer language models has already been demonstrated in Rydberg atom arrays on the square lattice Moss et al. (2023); Sprague and Czischek (2023); Czischek et al. (2022). RNNs possess two key properties that make them particularly well-suited for studying frustrated systems. Firstly, their ability to perform exact sampling helps mitigate frustration-induced ergodicity issues in quantum Monte Carlo. Secondly, the ability to define them in any spatial dimension without incurring additional computational intractability helps address challenges faced by techniques like DMRG, such as the increased computational cost stemming from increased entanglement in higher dimensions Hibat-Allah et al. (2020, 2022).
Our findings reveal that in the highly frustrated and highly entangled regimes of the system, the RNN predicts a paramagnetic phase without topological order, consistent with earlier QMC simulations Yan et al. (2023). However, in contrast to the QMC results in Ref. Yan et al., 2023, the RNN suggests the absence of a spin-glass phase. Nevertheless, in agreement with QMC, our numerical simulations indicate the emergence of a rugged optimization landscape, necessitating more optimization steps and thermal-like fluctuations to mitigate local minima in the RNN’s parameter landscape.
Overall, our results showcase the remarkable applicability and advantages of machine learning-based wave functions, particularly RNNs, in tackling challenging problems at the forefront of Rydberg atom array physics. These findings pave the way for further exploration of exotic phases and phenomena in highly frustrated quantum systems, harnessing the power of modern machine learning techniques to advance our understanding in this field.
II Methods
We focus our attention on an array of neutral atoms on the kagome lattice, interacting via laser excitation to atomic Rydberg states. We consider a lattice with periodic boundary conditions (PBC). The Hamiltonian of this system is given by Browaeys and Lahaye (2020); Samajdar et al. (2021):
[…]
$+\frac{1}{2}\sum_{i,j}V(||\bm{x}_{i}-\bm{x}_{j}||)\ket{r}_{i}\bra{r}_{i}\otimes\ket{r}_{j}\bra{r}_{j}.$
Here |$\ket{g}_{i},\ket{r}_{i}$ are respectively the ground and excited states of the Rydberg atom $i$. $\Omega$ is the Rabi frequency and $\delta$ is the laser detuning. $V(R)=C/R^{6}$ is the repulsive potential due to the dipole-dipole interaction between Rydberg atoms, which is responsible for the blockade mechanism Browaeys and Lahaye (2020). In practice, we define a blockade radius $R_{b}$ such that $V(R_{b}/a)=\Omega$, where $a$ is the distance between two neighbouring Rydberg atoms. Finally, we note that the sum over all possible pairs is truncated to a sum over neighbors separated by a distance cutoff $R_{c}=2$ or $R_{c}=4$. The choice $R_{c}=2$ is taken to compare with the DMRG results reported in Ref. Samajdar et al. (2021) as well as with the QMC findings in Ref. Yan et al. (2023).
II.1 Two dimensional RNNs
Figure 1: (a) An illustration of a positive RNN wave function. Each RNN cell receives an input $\bm{\sigma}_{n-1}$ and a hidden state $\bm{h}_{n-1}$ and outputs a new hidden state $\bm{h}_{n}$. This vector is taken as an input to the Softmax layer (denoted S) that computes the conditional probability $P_{i}$. (b) RNN autoregressive sampling scheme: after obtaining the probability vector $\bm{y}_{i}$ from the Softmax layer (S) in step $i$, we sample it to produce $\bm{\sigma}_{i}$. The latter is taken again as an input to the RNN along with the hidden state $\bm{h}_{i}$ to sample the following degree of freedom $\bm{\sigma}_{i+1}$. (c) Mapping of a Kagome lattice to a square lattice by embedding three atoms in a larger local Hilbert space. (d) A two-dimensional (2D) RNN with periodic boundary conditions for a $3\times 3$ lattice for illustration purposes. A bulk RNN cell receives two hidden states $\bm{h}_{i,j-1}$ and $\bm{h}_{i-(-1)^{j},j}$, as well as two input vectors $\bm{\sigma}_{i,j-1}$ and $\bm{\sigma}_{i-(-1)^{j},j}$ (not shown) as illustrated by the black solid arrows. RNN cells at the boundary receive additional hidden states $\bm{h}_{i,j+1}$ and $\bm{h}_{i+(-1)^{j},j}$, as well as two input vectors $\bm{\sigma}_{i,j+1}$ and $\bm{\sigma}_{i+(-1)^{j},j}$ (not shown), as demonstrated by the blue curved and solid arrows. The sampling path is taken as a zigzag path, as demonstrated by the dashed red arrows. The initial memory states of the 2D RNN and the initial inputs are null vectors, as indicated by the dashed black arrows.
The Rydberg Hamiltonian is stoquastic in nature Bravyi (2015), which implies that the ground-state wave function contains only positive amplitudes. This offers the opportunity to model the ground state with an RNN wave function with only positive amplitudes Hibat-Allah et al. (2020) which we adopt below. Complex extensions of RNN wave functions for non-stoquastic Hamiltonians have been explored in Refs. Hibat-Allah et al. (2020); Roth (2020); Hibat-Allah et al. (2022). To model a positive RNN wave function, we can express our ansatz in the computational basis as:
$\Psi_{\bm{\theta}}(\bm{\sigma})=\sqrt{p_{\bm{\theta}}(\bm{\sigma})},$
such that $\bm{\theta}$ corresponds to the variational parameters of the ansatz |$\ket{\Psi_{\bm{\theta}}}$, and $\bm{\sigma}=(\sigma_{1},\sigma_{2},\ldots,\sigma_{N})$ is a configuration of the Rydberg atoms. The main advantage of using RNN wave functions is the possibility of estimating observables through autoregressive sampling, which allows obtaining uncorrelated samples by construction Hibat-Allah et al. (2020). To do so, we model the joint probability $p_{\bm{\theta}}(\bm{\sigma})$ by constructing the conditionals $p_{\bm{\theta}}(\sigma_{i}|\sigma_{<i})$ by taking advantage of the probability chain rule
$p_{\bm{\theta}}(\bm{\sigma})=p_{\bm{\theta}}(\sigma_{1})p_{\bm{\theta}}(\sigma_{2}|\sigma_{1})\cdots p_{\bm{\theta}}(\sigma_{N}|\sigma_{N-1},\dots,\sigma_{2},\sigma_{1}).$
These conditional probabilities are obtained through a Softmax layer as follows:
$p_{\bm{\theta}}(\sigma_{i}|\sigma_{<i})=\bm{y}_{i}\cdot\bm{\sigma}_{i}.$
Here $\bm{y}_{i}=\text{Softmax}(U\bm{h}_{i}+\bm{c})$ where $U$ and $\bm{c}$ are, respectively, trainable weights and biases, and ‘Softmax’ corresponds to the normalizing Softmax activation function. Additionally, the memory (hidden) state $\bm{h}_{\rm i}$ is obtained recursively as Lipton et al. (2015):
$\bm{h}_{i}=f(W[\bm{h}_{i-1};\bm{\sigma}_{i-1}]+\bm{b}),$ (1)
such that [.;.][.;.][ . ; . ] is a concatenation of two vectors, while $\bm{\sigma}_{i-1}$ is a one-hot encoding of $\sigma_{i-1}$. These computations are illustrated in Fig. 1(a). $W$ and $\bm{b}$ are also trainable weights and biases, and $f$ is a user-defined activation function.
By virtue of the ‘Softmax’ activation function, the conditionals $p_{\bm{\theta}}(\sigma_{i}|\sigma_{<i})$ are normalized to one. This property implies that the RNN joint probability $p_{\bm{\theta}}$ is also normalized Hibat-Allah et al. (2020). Furthermore, by sampling the conditionals $p_{\bm{\theta}}(\sigma_{i}|\sigma_{<i})$ sequentially, as illustrated in Fig. 1(b), we can extract exact samples from the joint RNN probability $p_{\bm{\theta}}$. An attractive property of this scheme is the possibility to efficiently generate uncorrelated samples from different modes present in $p_{\bm{\theta}}$, whereas traditional Metropolis sampling scheme may get stuck in only one mode.
The atom configurations of a Rydberg atom array on a kagome lattice can be seen as an $L\times L\times 3$ array of binary degrees of freedom where $L$ is the size of each side of the lattice. As illustrated in Fig. 1(c), we can map our kagome lattice with a local Hilbert space of $2$ to a square lattice with an enlarged Hilbert space of size $82^{3}=$ 8 which we can study using our two-dimensional (2D) RNN wave function Hibat-Allah et al. (2023); Sprague and Czischek (2023).
To construct a 2D RNN ansatz that can handle PBC, we modify our RNN recursion in Eq. (1) to a two-dimensional recursion relation as:
$\bm{h}_{i,j}$ $=f\!\Big{(}W[\text{Neighbours}(\bm{h}_{i,j});\text{Neighbours}(\bm{\sigma}_{i,j})]+\bm{b}\Big{)}.$ (2)
$\bm{h}_{i,j}$ is a memory state with two indices for each atom in the two-dimensional lattice. Here ‘Neighbours($\bm{\sigma}_{i,j}$)’ returns a concatenation of the neighbours of $\bm{\sigma}_{i,j}$. The same observation goes for ‘Neighbours($\bm{h}_{i,j}$)’. These neighbours correspond to incoming vectors indicated by the black and blue arrows as illustrated in Fig. 1(d). More specifically, we define
$\text{Neighbours}(\bm{h}_{i,j})\equiv$ $\ [\bm{h}_{i-(-1)^{j},j};\bm{h}_{i,j-1};\bm{0};\bm{0}]$
on the bulk. On the boundaries, we take
$\text{Neighbours}(\bm{h}_{i,j})\equiv$ $\ [\bm{h}_{i-(-1)^{j},j};\bm{h}_{i,j-1};$
$\bm{h}_{i+(-1)^{j},j};\bm{h}_{i,j+1}].$
Note that PBC on the indices is assumed. The additional inputs $\bm{\sigma}_{i+(-1)^{j},j}$, $\bm{\sigma}_{i,j+1}$ and hidden states $\bm{h}_{i+(-1)^{j},j}$, $\bm{h}_{i,j+1}$ allow to take PBC into account and to introduce correlations between degrees of freedom at the boundaries. During the autoregressive sampling procedure, the input and hidden vectors are initialized to a null vector if not previously defined to preserve the autoregressive nature of our scheme, as illustrated in Fig. 1(b). Also, note that the particular choice of the indices is motivated by the zigzag sampling path. In this study, we use an advanced version of 2D RNNs incorporating the gating mechanism as previously done in Refs. Hibat-Allah et al. (2022); Casert et al. (2021); Luo et al. (2023). More details can be found in Appendix. A. Finally, since $\bm{h}_{i,j}$ is a summary of the history of the generated $\sigma_{<i,j}$, it is used to compute the conditional probabilities as follows:
$p_{\bm{\theta}}(\sigma_{i,j}|\sigma_{<i,j})=\text{Softmax}(U\bm{h}_{i,j}+\bm{c})\cdot\bm{\sigma}_{i,j}.$ (3)
II.2 Supplementing RNN optimization with annealing
To reach the ground state of the Rydberg atoms array Hamiltonian on the kagome lattice, we minimize the energy expectation value $E_{\bm{\theta}}=\bra{\Psi_{\bm{\theta}}}\hat{H}\ket{\Psi_{\bm{\theta}}}$ using the Variational Monte Carlo (VMC) scheme Becca and Sorella (2017) (see Appendix B). Due to the frustrated nature of the kagome lattice which can induce local minima in the VMC scheme, we leverage annealing with thermal-like fluctuations to mitigate local minima. This technique has been suggested and implemented in Refs. Roth (2020); Hibat-Allah et al. (2021, 2022); Roth et al. (2022); Khandoker et al. (2023); Hibat-Allah et al. (2023). In this case, we obtain a free-energy like cost function, defined as
$F_{\bm{\theta}}(n)=E_{\bm{\theta}}-T(n)S_{\rm classical}(p_{\bm{\theta}}),$ (4)
where $F_{\bm{\theta}}$ is a variational pseudo Free energy and $S_{\rm classical}$ is the classical Shannon entropy:
$S_{\rm classical}(p_{\bm{\theta}})=-\sum_{\bm{\sigma}}p_{\bm{\theta}}(\bm{\sigma})\log\left(p_{\bm{\theta}}(\bm{\sigma})\right).$ (5)
The previous sum goes over all classical Rydberg configurations $\{\bm{\sigma}\}$ in the computational $z$-basis. Note that $S_{\rm classical}$ is a pseudo-entropy that can be efficiently estimated using our RNN wave function as opposed to the quantum von Neumann entropy. Additionally, $T(n)$ is a pseudo-temperature that is annealed from some initial value $T_{0}$ to zero as follows: $T(n)=T_{0}(1-n/N_{\rm a})$ where $n\in[0,N_{\rm a}]$ and $N_{\rm a}$ is the total number of annealing steps. We present more details about the hyperparameters of our training scheme in Appendix. C.
II.3 Topological entanglement entropy
To investigate the existence of a topological property in the Rydberg atom arrays on the kagome lattice, we compute the topological entanglement entropy (TEE) Hamma et al. (2005a, b); Levin and Wen (2006); Kitaev and Preskill (2006); Flammia et al. (2009); Isakov et al. (2011); Wildeboer et al. (2017). For a gapped phase of matter, where the area law is satisfied, the Renyi-$2$ entanglement entropy follows the scaling law $S_{2}(A)=aL-\gamma+\mathcal{O}(L^{-1})$, assuming $A$ and $B$ is partition of the system, $L$ is the size of the boundary between $A$ and $B$ and $S_{2}(A)\equiv-\log(\text{Tr}(\rho^{2}_{A}))$. In this case, $\gamma$ is the so-called TEE. In this paper, we use the swap trick with our RNN wave function ansatz Hastings et al. (2010); Hibat-Allah et al. (2020); Wang and Davis (2020) to calculate the second Renyi entropy $S_{2}$ to extract the TEE $\gamma$.
We extract $\gamma$ using two different strategies, namely the Kitaev-Preskill construction Kitaev and Preskill (2006) and the Levin-Wen construction Levin and Wen (2006), illustrated in Fig. 2.
Figure 2: (a) A sketch of the parts $A$, $B$, and $C$ that we use for Kitaev-Preskill construction to compute the TEE. (b) Levin-Wen construction using the regions $A,B,C$ and $D$. For the Rydberg atoms Hamiltonian on a kagome lattice, each dot on the square lattice corresponds to a block of three binary degrees of freedom, as shown in Fig. 1(c).
The Kitaev-Preskill construction consists of choosing three subregions $A$, $B$, $C$ with geometries as shown in Fig. 2(a). The TEE can be then obtained by computing
$\gamma$ $=-S_{2}(A)-S_{2}(B)-S_{2}(C)+S_{2}(AB)$
$+S_{2}(AC)+S_{2}(BC)-S_{2}(ABC),$
where $S_{2}(A)$ is the second Renyi entropy of the subsystem $A$, and $AB$ is the union of $A$ and $B$ and similarly for the other terms. It is worth mentioning that finite size effects on $\gamma$ can be reduced by extrapolating the size of the subregions Kitaev and Preskill (2006); Furukawa and Misguich (2007). Finally, note that this approach combined with RNN wave functions was successful in extracting a non-zero TEE on the toric code and the hard-core Bose-Hubbard model on the kagome lattice Hibat-Allah et al. (2023).
The Levin-Wen construction allows to extract the TEE $\gamma$ by constructing four different subsystems $A_{1}=A\cup B\cup C\cup D,A_{2}=A\cup C\cup D,A_{3}=A\cup B\cup D$ and $A_{4}=A\cup D$ as illustrated in Fig. 2(b) such that Isakov et al. (2011):
$\gamma=\frac{-S_{2}(A_{1})+S_{2}(A_{2})+S_{2}(A_{3})-S_{2}(A_{4})}{2}.$
Note that finite size effects on $\gamma$ can be reduced by extrapolating the width and thickness of $A_{1},A_{2},A_{3}$ and $A_{4}$ Isakov et al. (2011); Furukawa and Misguich (2007).
Finally, we would like to highlight that our ability to study quantum systems with fully periodic boundary conditions is key to mitigating boundary effects, as opposed to cylinders used in DMRG Stoudenmire and White (2012); Gong et al. (2014), which can introduce boundary effects in the TEE value Samajdar et al. (2021).
III Results
Figure 3: In these panels, we focus on the Blockade radius $R_{b}=1.95$. (a) Plot of two point correlations $\langle n_{\bm{0}}n_{\bm{r}}\rangle$ with $\delta=3.3$ and for a system size $N=8\times 8\times 3$ and $R_{c}=2$. (b) Plots of the topological entanglement entropy versus $\delta$ for two different values of the cutoff radius $R_{c}$, using the Levin-Wen (LW) construction and the Kitaev-Preskill (KP) construction, for $N=8\times 8\times 3$. (c) A plot of the Edwards-Anderson order parameter $q_{\text{EA}}$ defined in Eq. (6) as a function of $\delta$ for $N=8\times 8\times 3$. (d) A plot of the density overlaps $\frac{1}{N}\sum_{i=1}^{N}n_{i}^{(1)}n_{i}^{(2)}$ and the spin overlaps $\frac{1}{N}\sum_{i=1}^{N}S_{i}^{(1)}S_{i}^{(2)}$ at $\delta=3.3$. Here $S_{i}=2n_{i}-1$, and (1) and (2) are labels for two sets of samples obtained from our optimized RNN, that are aggregated from $10$ different training seeds, for $N=6\times 6\times 3$ and $R_{c}=2$. For each seed, we generate $2\times 10^{5}$ independent samples and divide them into two sets.
According to the RNN numerics, our results show that the ground state at $R_{b}=1.95$ and $\delta=3.3$, which is suggested to be in the spin-liquid phase according to Ref. Samajdar et al. (2021), is rather a disordered state with no topological order. We first plot the correlations $\langle n_{\bm{0}}n_{\bm{r}}\rangle$ in Fig. 3(a). The results indicate that the extracted state has short-range correlations. To confirm the correctness of our variational implementation, we perform a sanity check and compare our ground state energies with QMC and DMRG as shown in Appendix D. We found a good agreement between our RNN energies and QMC as well as DMRG energies. Most importantly, we observe that our RNN results using only $d_{h}=100$ are more accurate compared to DMRG with a bond dimension $\chi=1000$ in the highly entangled regime at $R_{b}=1.95$ and $\delta=3.3$.
To investigate the existence of a spin liquid in this regime, we calculate the TEE $\gamma$ using the Kitaev-Preskill construction Kitaev and Preskill (2006) for a system size $L=8$ (see Fig. 2(a)), and for different values of $\delta\in[2.0,3.7]$ and $R_{c}=2,4$ at $R_{b}=1.95$. We also do the same using the Levin-Wen construction Isakov et al. (2011) in Fig. 2(b). Our results, illustrated in Fig. 3(b) suggest that the TEE extracted by the RNN is consistent with zero and different from $\ln(2)$ within error bars. These results suggest the non-existence of a spin liquid within our settings and also suggest that the state we find in this regime is a disordered state. Our findings are further corroborated by a recent QMC study Yan et al. (2023) and also by previous results in the literature suggesting that the paramagnetic ‘liquid’ phase in Ising systems on the kagome lattice is not exotic Nikolić and Senthil (2005); Moessner and Sondhi (2001); Moessner et al. (2000).
In this QMC study Yan et al. (2023), it was suggested that the region, around $R_{b}=1.95$ and the values of $\delta$ used in our study, contains an emergent spin-glass phase instead of a paramagnetic state. To verify this claim, we compute the Edwards-Anderson (EA) order parameters Edwards and Anderson (1975); Richards (1984), defined as:
$q_{\text{EA}}=\frac{\sum_{i=1}^{N}\langle n_{i}-\rho\rangle^{2}}{N\rho(1-\rho)},$ (6)
where $N$ is the system size, $n_{i}$ is the occupation number of site $i$ and $\rho=(\sum_{i=1}^{N}n_{i})/N$. Deviations of this order parameter from zero values are signals of the existence of a spin-glass phase. In Fig. 3(c), we plot this order parameter as a function of $\delta$ with $R_{c}=2,4$ and $R_{b}=1.95$. We find that the values of the order parameter are consistent with zero, as opposed to the results of QMC in Ref. Yan et al. (2023). Furthermore, we report in Fig. 3(d) the density-density overlap $\frac{1}{N}\sum_{i=1}^{N}n_{i}^{(1)}n_{i}^{(2)}$ and the spin-spin overlap $\frac{1}{N}\sum_{i=1}^{N}S_{i}^{(1)}S_{i}^{(2)}$ between different RNN samples at $R_{b}=1.95,\delta=3.3$, and $R_{c}=2$. Here labels (1) and (2) correspond to two independent sets of samples, which are obtained from optimized RNNs with $10$ different training seeds. The Gaussian nature of the overlap distribution in both representations is another indicator that there is no static signature of a spin-glass order Castellani and Cavagna (2005).
The discrepancy in our results and previous QMC findings Yan et al. (2023) could be related to emergent glassy dynamics in the QMC simulations, which results in very long auto-correlations times and thus in a non-ergodic behavior. To corroborate our findings, we run QMC simulations Merali et al. (2023) for larger inverse temperatures compared to Ref. Yan et al. (2023), namely for $\beta\geq 200$ and using $2.2\times 10^{6}$ Monte Carlo samples. We find that the QMC prediction for the EA order parameter is given as $q^{\text{QMC}}_{\text{EA}}=0.0000018(5)$ for $R_{b}=1.95,\delta=3.3$, a system size $8\times 8\times 3$, and for a radius cutoff $R_{c}=2$. The previous result agrees very well with our RNN findings in Fig. 3(c). This result is also confirmed by the good agreement between the RNN energies and the QMC energies as shown in Appendix. D. Our findings are further supported by the results of Ref. Yan et al. (2022), which suggests the possibility of transition in a quantum dimer model between nematic to paramagnetic to staggered states. In conclusion, our numerical investigation suggests that the long auto-correlation time could be a limiting factor in the QMC results reported in Ref. Yan et al. (2023).
We note that the emergence of a long autocorrelation time in QMC coincides with the emergence of a rugged optimization landscape, which in our simulations implies a longer number of annealing steps in our RNN simulations to achieve convergence. To demonstrate this point, we compute the structure factor
$S(\bm{q})=\frac{1}{N}\sum_{i,j}\langle n_{i}n_{j}\rangle\ e^{\text{i}\bm{q}\cdot(\bm{x}_{i}-\bm{x}_{j})}$ (7)
to extract the nature of the states obtained by our RNN ansatz and investigate their dependence on the number of annealing steps $N_{a}$. We expect the optimization landscape to be rougher as $N_{a}$ required to converge increases. Figs. 4(a-d) at $R_{b}=1.95$ and $\delta=3.3$ show that the RNN finds different states for different numbers of annealing steps $N_{a}$, until it converges to a state without ordering peaks, i.e., the paramagnetic state. In contrast, the nematic state at $R_{b}=1.7,\delta=3.3$ can be reached without the need for annealing, as illustrated by the structure factors at different $N_{a}$ in Fig. 4(e-h). These observations suggest an emergent rugged optimization landscape when optimizing our ansatz in the highly entangled regime. Finally, to find the optimal number of annealing steps $N_{a}$ in the highly entangled regime, we note that we conduct a scaling study as shown in Appendix E.
Figure 4: Plots of the structure factor for different values of the number of annealing steps $N_{a}$ at $R_{b}=1.95,\delta=3.3$ in panels (a-d) and at $R_{b}=1.7,\delta=3.3$ in panels (e-h). The color bars represent the magnitude of the structure factor $S(\bm{q})$. Additionally, these results correspond to a system size $N=6\times 6\times 3$. Note that we observe convergence to a paramagnetic state without ordering peaks beyond $N_{a}=1000$ for $R_{b}=1.95,\delta=3.3$. For $R_{b}=1.7,\delta=3.3$, we find that the nematic state is not affected by the choice of $N_{a}$ and can be obtained without a need for annealing.
IV Conclusions and Outlooks
In this paper, we demonstrate a successful application of recurrent neural network (RNN) wave functions to the task of investigating topological order on Rydberg atom arrays on kagome lattice. We use these architectures to estimate the second Renyi entropies using the swap trick Hibat-Allah et al. (2020). The latter allows us to compute the TEEs using the Kitaev-Preskill Kitaev and Preskill (2006) and the Levin-Wen Levin and Wen (2006) constructions. Furthermore, with the possibility of handling periodic boundary conditions in RNNs, the boundary effects on the TEE are reduced compared to DMRG, which has challenges with boundary effects on cylinders Gong et al. (2014).
Our main finding, suggested by the two-dimensional RNN wave functions results, points out that Rydberg atom arrays on the kagome lattice do not establish a $Z_{2}$ spin liquid in the highly entangled regime. This observation is also consistent with previous QMC studies Yan et al. (2023). Our RNN numerics also suggest that the highly entangled region corresponds to a trivial paramagnetic state and that there is no signature for spin glass order as opposed to the observations outlined in Ref. Yan et al. (2023). We believe that the ability of RNNs to generate uncorrelated samples from a multimodal distribution is a crucial factor in ruling out the spin-glass phase. Furthermore, supplementing RNNs with annealing turns out to be a valuable tool for mitigating local minima induced by the frustrated nature of the kagome lattice in the highly entangled regime. Additionally, we conclude that autocorrelation could be the main factor behind the observed spin glass phase observed in previous QMC simulations Yan et al. (2023).
Finally, we note that our method can be generalized to study other systems with potential topological order, such as the Rydberg atom arrays on the Ruby lattice Verresen et al. (2021); Semeghini et al. (2021); Giudici et al. (2022). One could also use quantum state tomography with RNNs Carrasquilla et al. (2019) in a wide variety of quantum simulators and also combine data with VMC to improve the variational results Bennewitz et al. (2021); Czischek et al. (2022); Moss et al. (2023). We also believe in the potential of RNN wave functions ansätzes in the discovery of new phases of matter with topological order. Overall, these results highlight the promising future of RNN wave functions Hibat-Allah et al. (2020); Roth (2020), language-model based wave functions, and neural quantum states Carleo and Troyer (2017) for investigating open questions and discovering new physics within the condensed matter community and beyond.
Code Availability
Our code is made publicly available at “http://github.com/mhibatallah/RNNWavefunctions”. The hyperparameters we use are given in Appendix. C.
Acknowledgments
We would like to thank Subir Sachdev, Anders Sandvik, and Arun Paramekanti for their helpful and inspiring discussions. Our RNN implementation is based on Tensorflow Abadi et al. (2015) and NumPy Harris et al. (2020). Computer simulations were made possible thanks to the Vector Institute computing cluster and the Digital Research Alliance of Canada cluster. We acknowledge support from Natural Sciences and Engineering Research Council of Canada (NSERC), the Shared Hierarchical Academic Research Computing Network (SHARCNET), Compute Canada, and the Canadian Institute for Advanced Research (CIFAR) AI chair program. This work is not related to the research being performed at AWS. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).
Appendix A Two dimensional periodic gated RNNs
In this Appendix, we share more details about our 2D gated RNN wave function implementation for periodic systems, which is used in this study to target the ground states of the Rydberg atom arrays on the kagome lattice. If we define
$\bm{h^{\prime}}_{i,j}$ $=[\bm{h}_{i-(-1)^{j},j};\bm{h}_{i,j-1};\bm{h}_{i+(-1)^{j},j};\bm{h}_{i,j+1}],$
$\bm{\sigma^{\prime}}_{i,j}$ $=[\bm{\sigma}_{i-(-1)^{j},j};\bm{\sigma}_{i,j-1};\bm{\sigma}_{i+(-1)^{j},j};\bm{\sigma}_{i,j+1}],$
then our gated 2D RNN wave function ansatz is based on the following recursion relations:
$\bm{h}_{i,j}$ $=\tanh\!\Big{(}W[\bm{\sigma^{\prime}}_{i,j};\bm{h^{\prime}}_{i,j}]+\bm{b}\Big{)},$
$\bm{u}_{i,j}$ $=\text{sigmoid}\!\Big{(}W_{g}[\bm{\sigma^{\prime}}_{i,j};\bm{h^{\prime}}_{i,j}]+\bm{b}_{g}\Big{)},$
$\bm{h}_{i,j}$ $=\bm{u}_{i,j}\odot\bm{\tilde{h}}_{i,j}+(1-\bm{u}_{i,j})\odot(U_{g}\bm{h^{\prime}}_{i,j}).$
A hidden state $\bm{h}_{i,j}$ can be obtained by combining a candidate state $\bm{\tilde{h}}_{i,j}$ and the neighbouring hidden states $\bm{h}_{i-1,j},\bm{h}_{i,j-1},\bm{h}_{i+1,j},\bm{h}_{i,j+1}$. The update gate $\bm{u}_{i,j}$ determines how much of the candidate hidden state $\bm{\tilde{h}}_{i,j}$ will be taken into account and how much of the neighboring states will be considered. With this combination, it is possible to mitigate some limitations of the vanishing gradient problems Zhou et al. (2016); Shen (2019). The weight matrices $W,W_{g},U_{g}$ and the biases $b,b_{g}$ are variational parameters of our RNN ansatz in addition to the Softmax layer parameters in Eq. (3). Note that we choose the size of the hidden state $\bm{h}_{i,j}$, which we denote as $d_{h}$, before optimizing our ansatz parameters. We note that the choice of the gated 2DRNN is motivated by its superiority compared to the non-gated 2DRNN on the task of finding the ground state of the 2D Heisenberg model Hibat-Allah et al. (2022).
Since we use an enlarged local Hilbert space with three atoms at each recursion step, the size of the Softmax layer output is defined as $2^{3}$. Additionally, each input $\bm{\sigma}_{i,j}$ is defined as a concatenation of the one-hot encoding of each of the three atoms. This means that $\bm{\sigma}_{i,j}$ is a six-dimensional vector.
Appendix B Variational monte carlo (VMC)
To optimize the energy expectation value of our RNN wave function $\Psi_{\bm{\theta}}$, we use the Variational Monte Carlo (VMC) scheme, which consists of using importance sampling to estimate the energy expectation value $E_{\bm{\theta}}=\bra{\Psi_{\bm{\theta}}}\hat{H}\ket{\Psi_{\bm{\theta}}}$ as follows Becca and Sorella (2017); Hibat-Allah et al. (2020):
$E_{\bm{\theta}}=\frac{1}{M}\sum_{i=1}^{M}E_{\text{loc}}(\bm{\sigma^{(i)}}),$
where the local energies $E_{\text{loc}}$ are defined as
$E_{\text{loc}}(\bm{\sigma})=\sum_{\bm{\sigma^{\prime}}}H_{\bm{\sigma}\bm{\sigma^{\prime}}}\frac{\Psi_{\bm{\theta}}(\bm{\sigma^{\prime}})}{\Psi_{\bm{\theta}}(\bm{\sigma})}.$
Here the configurations $\{\bm{\sigma^{(i)}}\}_{i=1}^{M}$ are sampled from our ansatz using autoregressive sampling. The choice of $M$ is a hyperparameter that can be tuned. Similarly, the gradients can be estimated as
$\partial_{\bm{\theta}}E_{\bm{\theta}}=\frac{1}{M}\sum_{i=1}^{M}\partial_{\bm{\theta}}\log\left(\Psi_{\bm{\theta}}^{*}(\bm{\sigma^{(i)}})\right)\left(E_{\text{loc}}(\bm{\sigma^{(i)}})-E_{\bm{\theta}}\right).$
Subtracting the mean energy $E_{\bm{\theta}}$ is helpful to achieve convergence as it reduces the variance of the gradients without biasing its expectation value Hibat-Allah et al. (2020, 2023). The gradient descent steps are performed using the Adam optimizer Kingma and Ba (2015). Similarly to the stochastic energy estimation, we can implement a similar procedure for the estimation of the variational pseudo-free energy $F_{\bm{\theta}}$ in Eq. (4). Ref. Hibat-Allah et al. (2021) provides more details in the supplementary information.
Appendix C Hyperparameters
For all models studied in this paper, we note that for each annealing step, we perform $N_{\rm train}=5$ gradient steps. Concerning the learning rate $\eta$, we choose $\eta=10^{-3}$ during the warm-up phase and the annealing phase and switch to a learning rate $\eta=10^{-4}$ in the convergence phase. To train RNN on the $8\times 8\times 3$ lattices, we pre-train using the optimized RNN on the $6\times 6\times 3$ lattice without using annealing, since the RNN is expected to start from a variational energy that is close to the ground state energy in the new system size.
In Tab. 1, we provide further details about the hyperparameters we choose in our study for the different models. For the estimation of the RNN energy, we use $2\times 10^{5}$ independent configurations. We also use $M=2\times 10^{6}$ independent samples for the estimation of the entanglement entropy along with their error bars. For the estimation of the TEE using Kitaev-Preskill or Levin-Wen constructions, we use the expression of the standard deviation of the sum of independent random variables to estimate the one standard deviation on $\gamma_{\rm RNN}$.
Figures Parameter Value
Fig. 3 ($N=6\times 6\times 3$) Number of memory units $d_{h}=100$
Number of training samples $M=500$
Initial pseudo-temperature $T_{0}=2$
Number of warm-up steps $N_{w}=1000$
Number of annealing steps $N_{a}=10000$
Number of convergence steps $N_{\text{conv}}=10000$
Number of samples for TEE estimation $M=10^{7}$
Number of samples for $q_{\text{EA}}$ estimation $M=2\times 10^{5}$
Fig. 3 ($N=8\times 8\times 3$) Number of memory units $d_{h}=100$
Number of training samples $M=500$
Initial pseudo-temperature $T_{0}=0$
Number of warm-up steps $N_{w}=0$
Number of annealing steps $N_{a}=0$ (pre-trained from $N=6\times 6\times 3$)
Number of training steps $N_{\text{train}}=10000$
Fig. 4 ($N=6\times 6\times 3$) Number of memory units $d_{h}=100$
Number of training samples $M=500$
Initial pseudo-temperature $T_{0}=2$
Number of warm-up steps $N_{w}=1000$
Number of convergence steps $N_{\text{conv}}=10000$
Number of sample for two-point correlations estimation $M=2\times 10^{5}$
Fig. 5 ($N=6\times 6\times 3$) Number of memory units $d_{h}=100$
Number of training samples $M=500$
Initial pseudo-temperature $T_{0}=2$
Number of warm-up steps $N_{w}=1000$
Number of convergence steps $N_{\text{conv}}=10000$
Number of samples for energy and density estimation $M=2\times 10^{5}$
Number of samples for $S_{2}$ estimation $M=2\times 10^{7}$
Table 1: A summary of the hyperparameters used to obtain the results reported in this paper.
Appendix D Numerical comparisons
In Tab. 2, we show a comparison between QMC’s and RNN’s energies per site for a system size $8\times 8\times 3$ and for a detuning $\delta=3.3$ and at the blockade radiuses $R_{b}=1.7,1.95$. These points correspond to the nematic and disordered phases, respectively. We note that our RNN-based ansatz provides energies with a relative error of less than $0.2\%$ compared to the QMC energies. The QMC simulations we run for Rydberg atom arrays are introduced in Ref. Merali et al. (2023). We use a finite-temperature QMC scheme run at several different values of $\beta$ until we observe convergence to the ground state. For each $\beta$, five independent simulations are taken and the convergence of observables is observed at $\beta=200$. Thus, to compute observables, we treat simulations with $\beta\geq 200$ as additional independent chains, giving us a total of 25 independent Markov chains at each parameter point. Each chain is allowed to warm-up for $10^{4}$ steps, after which $10^{6}$ sequential measurements were taken. With respect to the computation of the Edwards-Anderson order parameter, $q_{EA}$, we note that the analysis given in Ref. Yan et al. (2023) can give different results in the case of imperfect sampling. Ref. Yan et al. (2023) computes the order parameter independently for each Markov chain and then averages the results. This procedure can produce different results as each chain will only explore a subset of the QMC configuration space due to the presence of frustrated interactions. As a result, each chain’s estimate of the one-point function can be biased. Since $q_{EA}$ is a non-linear function of the one-point function, we must first aggregate the one-point functions generated by each Markov chain, and then compute $q_{EA}$. Lastly, to compute an estimate of the error in $q_{EA}$, we must account for auto-correlations and non-linearity simultaneously. This step is done by combining jackknife resampling with a binning procedure. To deal with auto-correlations, we first compute the one-point function on sequential “bins” of data; we found a bin size of $10^{4}$ to be sufficient, giving $100$ bins for each chain. Thus, we can consider each bin’s one-point function to be nearly uncorrelated, allowing us to directly apply the jackknife resampling procedure to these approximately independent bins.
To compare our RNN wave function ($d_{h}=100$) results with DMRG, we perform DMRG simulations using PastaQ Torlai and Fishman (2020) and ITensor Fishman et al. (2022) to further check the consistency of our RNN energies. In Tab. 3, we compare with DMRG using periodic boundary conditions. In the nematic phase $R_{b}=1.7,\delta=3.3$, we find an excellent match of the energies. For the disordered phase at $R_{b}=1.95,\delta=3.3$, our RNN energies are lower within an error of about $1\%$ and with orders of magnitude fewer parameters compared to DMRG. Furthermore, we choose the YC12 geometry used in Ref. Samajdar et al. (2021). We optimize our 2DRNN wave function ($d_{h}=60$) at $R_{b}=1.95$ and $\delta=3.5$. Our estimated energy is $-151.959(8)$ which is within 0.3% error compared to the DMRG energy provided in Ref. Samajdar et al. (2021).
QMC ($R_{c}=2$) 2DRNN ($R_{c}=2$) QMC ($R_{c}=4$) 2DRNN ($R_{c}=4$)
$R_{b}=1.7,\delta=3.3$0.79056(1) -0.790964(5) -0.77546(1) -0.775412(4)
$R_{b}=1.95,\delta=3.3$0.59785(1) -0.59657(1) -0.56445(1) -0.56401(3)
Table 2: A table of the energies per site obtained by QMC and 2DRNN for a system size $8\times 8\times 3$ with fully periodic boundary conditions and for different values of the cutoff radius $R_{c}$. The error bars in parentheses correspond to the one-standard deviation uncertainty in the QMC and RNN energy estimates.
Rydberg parameters DMRG ($R_{c}=2$) 2DRNN ($R_{c}=2$)
$R_{b}=1.7,\delta=3.3$0.790957 -0.790934(7)
$R_{b}=1.95,\delta=3.3$0.593828 -0.59700(4)
Table 3: A table comparing DMRG (with bond dimension $\chi=110$ for $R_{b}=1.7$ and $\chi=1000$ for $R_{b}=1.95$) and the 2DRNN ansatz in terms of the energies per site for a system size $6\times 6\times 3$ with fully periodic boundary conditions and for a cutoff radius $R_{c}=2$ and $\delta=3.3$. Note that a DMRG run on $R_{b}=1.95,\delta=3.3$ for $\chi=1500$ was not successful with a memory allocation of $120$ GB. To run our RNN simulations, an $80$ GB memory allocation was sufficient. The error bars in the parentheses correspond to the one-standard deviation uncertainty on the RNN energy.
Appendix E Annealing and local minima
In Fig. 5, we demonstrate the importance of incorporating annealing in the training of RNNs applied to Rydberg atom arrays on the kagome lattice. These experiments are carried out in the highly entangled regime at $R_{b}=1.95$ and $\delta=3.3$. In panel (a), we observe that the ground state energy improves with more annealing steps $N_{a}$. We also highlight that the density saturates close to $1/6$, which raises the possibility of an odd quantum spin liquid Yan et al. (2023); Samajdar et al. (2021). However, it is not a sufficient condition as this phase is found to be the paramagnetic state according to our RNN numerics. We also outline a saturation of the second Renyi entropy to a large value in the asymptotic limit of $N_{a}$. All these numerics suggest that $N_{a}=10000$ is a good choice to converge our RNN training.
Figure 5: Plots of the energy density $\langle\hat{H}\rangle/N$, density $\rho=(\sum_{i=1}^{N}n_{i})/N$ and the second Renyi entropy $S_{2}$ as function of the number of annealing steps $N_{a}$ for a system size $N=6\times 6\times 3$ and $R_{b}=1.95$ and $\delta=3.3$. We observe a saturation of all these observables in the asymptotic limit, which justifies the use of $N_{a}=10000$ in our numerical simulations in Fig. 3.